A Hole with a Perpendicular Tolerance of 0.010 and a Positional Tolerance of 0.030 is located at a distance of 0.010 from the designed center location. If the hole is drilled at an angle of 0.5 degrees from perpendicular, what is the maximum thickness of the material where this hole will be in specification?
Answer #1

Given that

Perpendicular tolerance = 0.010

Given that the hole is drilled 0.5 degrees off.

Given that the hole is 0.010 off position from its designed center.

we have to find the thickness of material t,

given that the hole is off by 0.010

we have the position tolerance 0.030, considering the same is the circle on which the deviated center of hole must be in.

the allowed offset of hole in radial direction will be half of the same i.e. 0.015

the hole center is already off by 0.01,

so the available tolerance for perpendicular deviation is 0.015 - 0.010 = 0.0050

from the figure for 0.5 degrees of deviation.

sin = opposite side / hypo = 0.010 / hyp

sin 0.5 = 0.010/ hyp

hyp = 0.010/ Sin 0.5 = 1.459

and

cos 0.5 = adj side/ hyp

adj side = cos 0.5 * hyp

substituting the value of hyp

adj side = cos 0.5 * 1.459 = 1.4589 , the maximum thickness of material possible for current tolerance criteria.

Thanks......... please don't forget to like the solution.